# TIPS Over Tricks: Simple Prompts for Effective Zero-Shot Anomaly Detection

### v1

@@ Abstract
Anomaly detection identifies departures from expected behavior in safety-critical settings. When target-domain normal data are unavailable, zero-shot anomaly detection (ZSAD) leverages vision–language models (VLMs). However, CLIP’s coarse image–text alignment limits both localization and detection due to (i) spatial misalignment and (ii) weak sensitivity to fine-grained anomalies; prior work compensates with complex auxiliary modules yet largely overlooks the choice of backbone. We revisit the backbone and use TIPS—a VLM trained with spatially aware objectives. While TIPS alleviates CLIP’s issues, it exposes a distributional gap between global and local features. We address this with decoupled prompts—fixed for image-level detection and learnable for pixel-level localization—and by injecting local evidence into the global score. Without CLIP-specific tricks, our TIPS-based pipeline improves image-level performance by 1.1–3.9% and pixel-level by 1.5–6.9% across seven industrial datasets, delivering strong generalization with a lean architecture. Code is available at github.com/AlirezaSalehy/Tipsomaly.
@@ 1 Introduction
To mitigate these issues, prior ZSAD methods introduce non-learnable refinements (“VV” attention [34, 6]) and learnable modifications to the vision encoder (trainable visual prompts [4], feature adapters [5]). Although helpful, these adaptations cannot fully restore the spatial alignment and patch–text correspondence absent in CLIP inherently, and increase architectural complexity. Moreover, previous analyzes suggest that learnable adaptations can overfit train data, weakening the cross-domain generalization required for zero-shot anomaly detection [26]. For example, AdaCLIP’s [4] joint tuning of text and vision encoders can improve image-level detection, but transfers poorly to pixel-level localization, underscoring the limits of CLIP-centric adaptations.
In this paper, we revisit an underexplored lever for ZSAD—the backbone—and investigate TIPS’s [19] capabilities, a more spatially aware model for the task, showing it sidesteps several CLIP limitations and enables a simpler pipeline without complex adaptations. However, TIPS is trained with a generic vision–language objective, making direct utilization suboptimal for anomaly detection. We evaluated learnable prompts and fixed textual templates and found a consistent distributional gap between local and global features: when learnable prompts are trained with an additional global objective (normal vs. anomalous classification) alongside the local loss, image-level performance improves but pixel-level alignment degrades, indicating a mismatch between local and global anomaly representations.
@@ 2.1 Problem Definition
thereafter the final Image- and pixel-level decisions are obtained by thresholding these scores.
@@ 2.4 Learnable Localization Prompts
Passing each element of $P_{l}$ through the text encoder $T(\cdot)$ yields the normal/abnormal localization prototypes $G_{l}=\{g^{n}_{l},\,g^{a}_{l}\}$ which used to score patch embeddings $Z^{\mathrm{M}}$ via similarity.
@@ 2.5 Training Objective & Inference
Having obtained the textual embeddings specialized for detection $G_{f}$ and localization $G_{l}$ and visual embeddings $e\in\{g_{o}^{i},g_{s}^{i},Z_{j,k}\}$, where $Z_{j,k}$ denotes $(j_{th},k_{th})$ patch embedding of unflattened local feature $Z^{M}$. for input image $x$, we can now compute the likelihood of each visual embedding belonging to the anomalous class ($p_{a}$) by applying the Softmax function to similarity scores:
@@
Train. We learn the normal/abnormal patch-level visual distribution via learnable prompts $P_{l}$, while keeping the vision encoder $\Phi_{v}$ and text encoder $\Phi_{t}$ frozen and employing local loss function, Focal [18] and Dice [29] loss. For output feature map $Z^{M}$, we compute normal and anomaly maps $\hat{S}_{n}$ and $\hat{S}_{a}$ based on the equation 1, then apply bilinear upsampling $\text{Up}(\cdot)$ to match the anomaly mask $S\in\mathbb{R}^{H\times W}$:
@@
Inference. For each input image $x$, after computing visual outputs including local features $Z^{M}$ and global embeddings $g_{o}^{i}$ and $g_{s}^{i}$ alongside detection and localization textual embeddings $G_{f}$ and $P_{l}$ in forward pass, low-resolution anomaly maps $\hat{S}_{a}$ is calculated then bilinearly upsampled and smoothed with Gaussian filter to obtain final anomaly map $\hat{S}$:
@@ 3 Experiments
We evaluate on diverse 14 real-world anomaly detection datasets covering industrial defects and medical abnormalities, including MVTec-AD [2], VisA [35], DTD-Synthetic [1], SDD [30], BTAD [20], DAGM [33], and MPDD [14] for industrial and ISIC [10], CVC-ColonDB [31], CVC-ClinicDB [3], TN3K [8], BrainMRI [15], HeadCT [16], and BR35H [11] for medical domain. Following prior works [34, 4], we train on MVTec and test generalization on the others; for MVTec evaluation, we use VisA. In both cases, the categories are distinct and non-overlapping with those in the other datasets. For metrics, we follow previous studies [5, 34], using AUROC, AP, and F1-max for image-level and AUROC, AUPRO, and F1-max for pixel-level, and report the mean over categories as the dataset-level score.
@@
Table 2: Ablation study key component. Performance is reported at image-level (AUROC, AP) and pixel-level (AUROC, AUPRO) on MVTec and VisA. Higher values indicate improved performance. Default configuration is highlighted.
@@
Implementation Details. We use the publicly available TIPS-L/14 HR [19] as our default backbone. It has 487.1M parameters, comparable to the CLIP-L/14@336 (428M) used in prior works. We use 8 learnable tokens for each of $T^{\mathrm{n}}$ and $T^{\mathrm{a}}$. For optimizer, we use Adam with a learning rate of 0.001 and betas=(0.5, 0.999), and train for 2 epochs with a batch size of 8. The random seed for all algorithms is set to 111 to ensure reproducibility. Input images to the vision tower are resized to 518×518 and undergo the same normalization during both training and inference. Experiments are conducted using PyTorch 2.5 on a NVIDIA A6000 GPU.
@@ 3.3 Ablation Studies
Decoupled Prompting. Table 2(a) compares fixed textual prompts, learnable prompts, and a decoupled design. Fixed prompts favor image-level classification but collapse on segmentation, whereas learnable prompts excel at pixel-level localization yet underperform in classification. Decoupling resolves this inconsistency: Using fixed prompts for image-level scoring and learnable prompts for pixel-level maps, retaining strong image-level and pixel-level performance.
@@
Image-level Strategies Table 2(c) assesses the discriminative ability of $g^{i}_{o}$ and $g^{i}_{s}$ for image-level anomaly detection, along with aggregation strategies (mean vs. max), with $p_{a}(g^{i}_{s},G_{f})$ yielding the best discrimination. Finally, augmenting the global score with the strongest local evidence, $\max(\hat{S}_{a})$, produces a net gain on the ten image-level datasets reported in Table 1, though this is not uniformly reflected by the two ablation sets.

### v2

@@ Abstract
Anomaly detection identifies departures from expected behavior in safety-critical settings. When target-domain normal data are unavailable, zero-shot anomaly detection (ZSAD) leverages vision–language models (VLMs). However, CLIP’s coarse image–text alignment limits both localization and detection due to (i) spatial misalignment and (ii) weak sensitivity to fine-grained anomalies; prior works compensate with complex auxiliary modules yet largely overlook the choice of backbone. We revisit the backbone and use TIPS—a VLM trained with spatially aware objectives. While TIPS alleviates CLIP’s issues, it exposes a distributional gap between global and local features. We address this with decoupled prompts—fixed for image-level detection and learnable for pixel-level localization—and by injecting local evidence into the global score. Without CLIP-specific tricks, our TIPS-based pipeline improves image-level performance by 1.1–3.9% and pixel-level by 1.5–6.9% across seven industrial datasets, delivering strong generalization with a lean architecture. Code is available at github.com/AlirezaSalehy/Tipsomaly.
@@ 1 Introduction
To mitigate these issues, prior ZSAD methods introduce non-learnable refinements (“VV” attention [34, 6]) and learnable modifications to the vision encoder (trainable visual prompts [4], feature adapters [5]). Although helpful, these adaptations cannot fully restore the spatial alignment and patch–text correspondence absent in CLIP inherently, and increase architectural complexity. Moreover, previous analyses suggest that learnable adaptations can overfit train data, weakening the cross-domain generalization required for zero-shot anomaly detection [26]. For example, AdaCLIP’s [4] joint tuning of text and vision encoders can improve image-level detection, but transfers poorly to pixel-level localization, underscoring the limits of CLIP-centric adaptations.
In this paper, we revisit an underexplored lever for ZSAD—the backbone—and investigate TIPS’s [19] capabilities, a more spatially aware model for the task, showing it sidesteps several CLIP limitations and enables a simpler pipeline without complex adaptations. However, TIPS is trained with a general vision–language objective, making direct utilization suboptimal for anomaly detection. We evaluated learnable prompts and fixed textual templates and found a consistent distributional gap between local and global features: when learnable prompts are trained with an additional global objective (normal vs. anomalous classification) alongside the local loss, image-level performance improves but pixel-level alignment degrades, indicating a mismatch between local and global anomaly representations.
@@ 2.1 Problem Definition
Thereafter, the final image- and pixel-level decisions are obtained by thresholding these scores.
@@ 2.4 Learnable Localization Prompts
Passing each element of $P_{l}$ through the text encoder $T(\cdot)$ yields the normal/abnormal localization prototypes $G_{l}=\{g^{n}_{l},\,g^{a}_{l}\}$ which are used to score patch embeddings $Z^{\mathrm{M}}$ via similarity.
@@ 2.5 Training Objective & Inference
Having obtained the textual embeddings specialized for detection $G_{f}$ and localization $G_{l}$, as well as the visual embeddings $e\in\{g_{o}^{i},g_{s}^{i},Z_{j,k}\}$, where $Z_{j,k}$ denotes the $(j,k)$-th patch embedding in the unflattened local feature map $Z^{M}$, we compute the probability that each visual embedding belongs to the anomalous class, denoted by $p_{a}$, by applying the Softmax function to the corresponding similarity scores:
@@
Training. We learn the normal/abnormal patch-level visual distribution via learnable prompts $P_{l}$, while keeping the vision encoder $\Phi_{v}$ and text encoder $\Phi_{t}$ frozen and employing local loss functions: Focal loss [18] and Dice loss [29]. For output feature map $Z^{M}$, we compute normal and anomaly maps $\hat{S}_{n}$ and $\hat{S}_{a}$ based on Eq. (1), then apply bilinear upsampling $\text{Up}(\cdot)$ to match the anomaly mask $S\in\mathbb{R}^{H\times W}$:
@@
Inference. For each input image $x$, after computing visual outputs including local features $Z^{M}$ and global embeddings $g_{o}^{i}$ and $g_{s}^{i}$ alongside detection and localization textual embeddings $G_{f}$ and $G_{l}$ in forward pass, low-resolution anomaly map $\hat{S}_{a}$ is calculated, then bilinearly upsampled and smoothed with Gaussian filter to obtain final anomaly map $\hat{S}$:
@@ 3 Experiments
We evaluate on 14 diverse real-world anomaly detection datasets covering industrial defects and medical abnormalities, including MVTec-AD [2], VisA [35], DTD-Synthetic [1], SDD [30], BTAD [20], DAGM [33], and MPDD [14] for industrial and ISIC [10], CVC-ColonDB [31], CVC-ClinicDB [3], TN3K [8], BrainMRI [15], HeadCT [16], and BR35H [11] for medical domain. Following prior works [34, 4], we train on MVTec and test generalization on the others; for MVTec evaluation, we use VisA. In both cases, the categories are distinct and non-overlapping with those in the other datasets. For metrics, we follow previous studies [5, 34], using AUROC, AP, and F1-max for image-level and AUROC, AUPRO, and F1-max for pixel-level, and report the mean over categories as the dataset-level score.
@@
Table 2: Ablation study of key components. Performance is reported at image-level (AUROC, AP) and pixel-level (AUROC, AUPRO) on MVTec and VisA. Higher values indicate improved performance. Default configuration is highlighted.
@@
Implementation Details. We use the publicly available TIPS-L/14 HR [19] as our default backbone. It has 487.1M parameters, comparable to the CLIP-L/14@336 (428M) used in prior works. We use 8 learnable tokens for each of $T^{\mathrm{n}}$ and $T^{\mathrm{a}}$. For optimizer, we use Adam with a learning rate of 0.001 and betas=(0.5, 0.999), and train for 2 epochs with a batch size of 8. The random seed for all algorithms is set to 111 to ensure reproducibility. Input images to the vision tower are resized to 518×518 and undergo the same normalization during both training and inference. Experiments are conducted using PyTorch 2.5 on an NVIDIA A6000 GPU.
@@ 3.3 Ablation Studies
Decoupled Prompting. Table 2(a) compares fixed textual prompts, learnable prompts, and a decoupled design. Fixed prompts favor image-level classification but collapse on segmentation, whereas learnable prompts excel at pixel-level localization yet underperform in classification. Decoupling resolves this inconsistency: we use fixed prompts for image-level scoring and learnable prompts for pixel-level maps, retaining strong performance at both levels.
@@
Image-level Strategies. Table 2(c) assesses the discriminative ability of $g^{i}_{o}$ and $g^{i}_{s}$ for image-level anomaly detection, along with aggregation strategies (mean vs. max), with $p_{a}(g^{i}_{s},G_{f})$ yielding the best discrimination. Finally, augmenting the global score with the strongest local evidence, $\max(\hat{S}_{a})$, produces a net gain on the ten image-level datasets reported in Table 1, though this is not uniformly reflected by the two ablation sets.